OpenStudy (anonymous):

if 250 mg of radioactive element decays to 200mg in 48 hours, find the half-life of the element

OpenStudy (anonymous):

any particular decay formula?

OpenStudy (anonymous):

I dont know thats why I am asking...

OpenStudy (anonymous):

m(t)=m e^(-rt) m(t) = amount after t hours m = initial amount r = rate t = time imma use this then first we have to find the rate of decay, r, by substitute the given information 200 = 250 e^(-r48) 4/5 = e^(-r48) ln(4/5) = -r48 ln(4/5)/(-48) = r still with me? :d

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

r = 0.0046 for the half life, it means m(t) = 1/2 m. so~ 1 = 2 e^(-0.0046t) 1/2 = e^(-0.0046t) ln(1/2) = -0.0046t ln(1/2)/(-0.0046) = t

OpenStudy (anonymous):

okayy

OpenStudy (anonymous):

got it all? o.o"

OpenStudy (anonymous):

yeah thank you